# Learning Language-guided Adaptive Hyper-modality Representation for Multimodal Sentiment Analysis

### v1

@@ 1 Introduction
Most recent MSA methods can be grouped into two categories: representation learning-centered methods (Hazarika et al., 2020; Yang et al., 2022; Yu et al., 2021; Han et al., 2021; Guo et al., 2022) and multimodal fusion-centered methods (Zadeh et al., 2017; Liu et al., 2018; Tsai et al., 2019a; Huang et al., 2020). The representation learning-centered methods mainly focus on learning refined modality semantics that contains rich and varied human sentiment clues, which can further improve the efficiency of multimodal fusion for relationship modelling. On the other hand, the multimodal fusion-centered methods mainly focus on directly designing sophisticated fusion mechanisms to obtain a joint representation of multimodal data. In addition, some works and corresponding ablation studies (Hazarika et al., 2020; Rahman et al., 2020; Guo et al., 2022) further imply that various modalities contribute differently to recognition, where language modality stands out as the dominant one. We note, however, information from different modalities may be ambiguous and conflicting due to sentiment-irrelevance, especially from non-dominating modalities (e.g., lighting and head pose in video and background noise in audio). Such disruptive information can greatly limit the performance of MSA methods. We have observed this phenomenon in several public datasets (see Section 4.5.1) and an illustration is in Figure 1. To the best of our knowledge, there has never been prior work explicitly and actively taking this factor into account. Motivated by the above observation, we propose a novel Adaptive Language-guided Multimodal Transformer (ALMT) to improve the performance of MSA by addressing the adverse effects of disruptive information in visual and audio modalities. In ALMT, each modality is first transformed into a unified form by using a Transformer with initialized tokens. This operation not only suppresses the redundant information across modalities, but also compresses the length of long sequences to facilitate efficient model computation. Then, we introduce an Adaptive Hyper-modality Learning (AHL) module that uses different scales of language features with dominance to guide the visual and audio modalities to produce the intermediate hyper-modality token, which contains less sentiment-irrelevant information. Finally, we apply a cross-modality fusion Transformer with language features serving as query and hyper-modality features serving as key and value. In this sense, the complementary relations between language and visual and audio modalities are implicitly reasoned, achieving robust and accurate sentiment predictions. In summary, the major contributions of our work can be summarized as:
@@
  • ALMT achieves state-of-the-art performance in several public and widely adopted MSA datasets. We further provide in-depth analysis with rich empirical results to demonstrate the validity and necessity of the proposed approach.
Figure 2: Processing pipeline of the proposed ALMT for multimodal sentiment analysis (MSA). With the multimodal input, we first apply three Transformer layers to embed modality features with low redundancy. Then, we employ a Hyper-modality Learning (AHL) module to learn a hyper-modality representation from visual and audio modalities under the guidance of language features at different scales. Finally, a Cross-modality Fusion Transformer is applied to incorporate hyper-modality features based on their relations to the language features, thus obtaining a complementary and joint representation for MSA
Figure 3: An example of the Adaptive Hyper-modality Learning (AHL) Layer
@@ 2.1 Multimodal Sentiment Analysis
As mentioned in the section above, most previous MSA methods are mainly classified into two categories: representation learning-centered methods and multimodal fusion-centered methods. For representation learning-centered methods, Hazarika et al. (2020) and Yang et al. (2022) argued representation learning of multiple modalities as a domain adaptation task. They respectively used metric learning and adversarial learning to learn the modality-invariant and modality-specific subspaces for multimodal fusion, achieving advanced performance in several popular datasets. Han et al. (2021) proposed a framework named MMIM that improves multimodal fusion with hierarchical mutual information maximization. Rahman et al. (2020) and Guo et al. (2022) devised different architectures to enhance language representation by incorporating multimodal interactions between language and non-verbal behavior information. However, these methods do not pay enough attention to sentiment-irrelevant redundant information that is more likely to be present in visual and audio modalities, which limits the performance of MSA. For multimodal fusion-centered methods, Zadeh et al. (2017) proposed a fusion method (TFN) using a tensor fusion network to model the relationships between different modalities by computing the cartesian product. Tsai et al. (2019a) and Huang et al. (2020) introduced a multimodal Transformer to align the sequences and model long-range dependencies between elements across modalities. However, these methods directly fuse information from uni-modalities, which is more accessible to the introduction of sentiment-irrelevant information, thus obtaining sub-optimal results.
@@ 3.3 Modality Embedding
$H^{1}_{m}=\operatorname{E^{0}_{m}}(\operatorname{concat}(H^{0}_{m},U_{m}),\theta_{E^{0}_{m}})\in\mathbb{R}^{T\times d},$ (1)
where $H^{1}_{m}$ is the unified feature of each modality $m$ with a size of $T\times d$, $\operatorname{E^{0}_{m}}$ and $\theta_{E^{0}_{m}}$ respectively represent the modality feature extractor and corresponding parameters, $\operatorname{concat}(\cdot)$ represent the concatenation operation. In practice, $T$ and $d$ are set to 8 and 128, respectively. The structure of the transformer layer is designed as the same as the Vision Transformer (VIT) (Dosovitskiy et al., 2021) with a depth setting of 1. Moreover, it is worth noting that transferring the essential modality information to initialized low-dimensional tokens is beneficial to decrease the redundant information that is irrelevant to human sentiment, thus achieving higher efficiency with lesser parameters.
@@ 3.4.1 Construction of Two-scale language Features
$H^{i}_{l}=\operatorname{E}^{i}_{l}(H^{i-1}_{l},\theta_{E_{l}^{i}})\in\mathbb{R}^{T\times d},$ (2)
@@ 3.4.2 Adaptive Hyper-modality Learning Layer
where $\operatorname{softmax}$ represents weight normalization operation, $W_{Q_{l}}\in\mathbb{R}^{d\times d_{k}}$ and $W_{K_{a}}\in\mathbb{R}^{d\times d_{k}}$ are learnable parameters, $d_{k}$ is the dimension of each attention head. In practice, we used 8-head attention and set $d_{k}$ to 16. Similar to $\alpha$, $\beta$ represents the similarity matrix between language modality and visual modality:
@@
where $W_{K_{v}}\in\mathbb{R}^{d\times d_{k}}$ is learnable. Then the hyper-modality features $H^{j}_{hyper}$ can be updated by weighted audio features and weighted visual features as:
@@ 3.6 Overall Learning Objectives
where $N_{b}$ is the number of samples in the training set, $y^{n}$ is the sentiment label of the $n$-th sample. $\hat{y}^{n}$ is the prediction of our ALMT. In addition, thanks to our simple optimization goal, compared with advanced methods Hazarika et al. (2020); Yu et al. (2021) with multiple optimization goals, ALMT is much easier to train without tuning extra hyper-parameters. More details are shown in section 4.5.4.
Table 1: Comparison on MOSI and MOSEI. Note: the best result is highlighted in bold; * represents the result is from Hazarika et al. (2020); $\dagger$ represents the result is from Mao et al. (2022) and its corresponding GitHub page4.1
@@
Table 2: Comparison results on CH-SIMS. Note: the best result is highlighted in bold; $\dagger$ represents the result is from Mao et al. (2022) and its corresponding GitHub page4.1
@@ 4.1 Datasets
We conducted extensive experiments on three popular datasets (i.e., MOSI Zadeh et al. (2016), MOSEI (Zadeh et al., 2018), and CH-SIMS (Yu et al., 2020)). MOSI. The dataset comprises 2,199 multimodal samples encompassing visual, audio, and language modalities. Specifically, the training set consists of 1,284 samples, the validation set contains 229 samples, and the test set encompasses 686 samples. Each individual sample is assigned a sentiment score ranging from -3 (indicating strongly negative) to 3 (indicating strongly positive). MOSEI. The dataset comprises 22,856 video clips collected from YouTube with a diverse factors (e.g., spontaneous expressions, head poses, occlusions, illuminations). This dataset has been categorized into 16,326 training instances, 1,871 validation instances, and 4,659 test instances. Each instance is meticulously labeled with a sentiment score ranging from -3 to 3. And the sentiment scores from -3 to 3 indicate most negative to most positive. CH-SIMS. It is a Chinese multimodal sentiment dataset that comprises 2,281 video clips collected from variuous sources, such as different movies and TV serials with spontaneous expressions, various head poses, etc. It is divided into 1,368 training samples, 456 validation samples, and 457 test samples. Each sample is manually annotated with a sentiment score from -1 (strongly negative) to 1 (strongly positive). 11footnotetext: https://github.com/thuiar/MMSA/blob/master/results/result-stat.md
@@ 4.4 Performance Comparison
Table 1 and Table 2 list the comparison results of our proposed method and state-of-the-art methods on the MOSI, MOSEI, and CH-SIMS, respectively. As shown in the Table 1, the proposed ALMT obtained state-of-the-art performance in almost all metrics. On the task of more difficult and fine-grained sentiment classification (Acc-7), our model achieves remarkable improvements. For example, on the MOSI dataset, ALMT achieved a relative improvement of 1.69% compared to the second-best result obtained by CHFN. It demonstrates that eliminating the redundancy of auxiliary modalities is essential for effective MSA. Moreover, it is worth noting that the scenarios in SIMS are more complex than MOSI and MOSEI. Therefore, it is more challenging to model the multimodal data. However, as shown in the Table 2, ALMT achieved state-of-the-art performance in all metrics compared to the sub-optimal approach. For example, compared to Self-MM, it achieved relative improvements with 1.44% on Acc-2 and 1.40% on the corresponding F1, respectively. Achieving such superior performance on SIMS with more complex scenarios demonstrates ALMT’s ability to extract effective sentiment information from various scenarios.
@@ 4.5.1 Effects of Different Modalities
To better understand the influence of each modality in the proposed ALMT, Table 3 reports the ablation results of the subtraction of each modality to the ALMT on the MOSI and CH-SIMS datasets, respectively. It is shown that, if the AHL is removed based on the subtraction of each modality, the performance decreases significantly in all metrics. This phenomenon demonstrates that AHL is beneficial in reducing the sentiment-irrelevant redundancy of visual and audio modalities, thus improving the robustness of MSA. In addition, we note that after removing the video and audio inputs, the performance of ALMT remains relatively high. Therefore, in the MSA task, we argue that eliminating the sentiment-irrelevant information that appears in auxiliary modalities (i.e., visual and audio modalities) and improving the contribution of auxiliary modalities in performance should be paid more attention to.
Table 3: Effects of different Modalities. Note: the best result is highlighted in bold
@@ 4.5.2 Effects of Different Components
Table 4: Effects of different components. Note: the best result is highlighted in bold
@@ 4.5.3 Effects of the Guidance of Different Language Features in AHL
To discuss the effect of the guidance of different language features in AHL, we show the ablation result of different guidance settings on MOSI and CH-SIMS in Table 5. In practice, we replace the AHL layer that do not require language guidance with MLP layer. Obviously, we can see that the ALMT can obtain the best performance when all scals of language features (i.e., $H_{l}^{1}$, $H_{l}^{2}$, $H_{l}^{3}$) involve the guidance of hyper-modality learning. In addition, we found that the model is more difficult to converge when AHL is removed. It indicates that sentiment-irrelevant and conflicting information visual and audio modalities may limit the improvement of the model.
Table 5: Effects of different Guidance of Different Language Features in AHL. Note: the best result is highlighted in bold
@@
4.5.4 Visualization of Convergence Performance
In Figure 4, we compared the convergence hehavior of ALMT with three state-of-the-art methods (i.e., MulT, MISA and Self-MM) on CH-SIMS. We choose the MAE curve for comparison as MAE indicates the model’s ability to predict fine-grained sentiment. Obviously, on the training set, although Self-MM converges the fastest, its MAE of convergence is larger than ALMT at the end of the epoch. On the validation set, ALMT seems more stable compared to other methods, while the curves of self-MM, MulT, and MISA show relatively more dramatic fluctuations. It demonstrates that the ALMT is easier to train and has a better generalization capability.
Figure 4: Visualization of convergence performance on the train and validation sets of CH-SIMS. (a) The comparison of MAE curves on the training set; (b) the comparison of MAE curves on the validation set. Note: the results of other methods reproduced by authors from open source code with default hyper-parameters
4.5.5 Visualization of Different Representations
Figure 5: Visualization of different representations in 3D space by using t-SNE
4.5.6 Visualization of Attention in AHL
Figure 6: Visualization of average attention weights from the last AHL layer on CH-SIMS dataset. (a) Average attention matrix $\alpha$ between language and audio modalities; (b) average attention matrix $\beta$ between language and visual modalities. Note: darker colors indicate higher attention weights for learning
5 Conclusion
In this paper, a novel Adaptive Language-guided Multimodal Transformer (ALMT) is proposed to better model sentiment cues for robust Multimodal Sentiment Analysis (MSA). ALMT consists of three major components, i.e., modality embedding, adaptive hyper-modality learning, and multimodal fusion. Due to effectively suppressing the adverse effects of redundant information in visual and audio modalities, the proposed method achieved highly improved performance on several popular datasets. We further present rich in-depth studies investigating the reasons behind the effectiveness, which may potentially advise other researchers to better handle MSA-related tasks.
Limitations
Appendix A Effects of different Query, Key, and Value Settings in Fusion Transformer
Table 6: Effect of different Query, Key, and Value Setting in Fusion Transformer
Q K & V MOSI CH-SIMS

### v2

@@ 1 Introduction
Most recent MSA methods can be grouped into two categories: representation learning-centered methods (Hazarika et al., 2020; Yang et al., 2022; Yu et al., 2021; Han et al., 2021; Guo et al., 2022) and multimodal fusion-centered methods (Zadeh et al., 2017; Liu et al., 2018; Tsai et al., 2019a; Huang et al., 2020). The representation learning-centered methods mainly focus on learning refined modality semantics that contains rich and varied human sentiment clues, which can further improve the efficiency of multimodal fusion for relationship modelling. On the other hand, the multimodal fusion-centered methods mainly focus on directly designing sophisticated fusion mechanisms to obtain a joint representation of multimodal data. In addition, some works and corresponding ablation studies (Hazarika et al., 2020; Rahman et al., 2020; Guo et al., 2022) further imply that various modalities contribute differently to recognition, where language modality stands out as the dominant one. We note, however, information from different modalities may be ambiguous and conflicting due to sentiment-irrelevance, especially from non-dominating modalities (e.g., lighting and head pose in video and background noise in audio). Such disruptive information can greatly limit the performance of MSA methods. We have observed this phenomenon in several datasets (see Section 4.5.1) and an illustration is in Figure 1. To the best of our knowledge, there has never been prior work explicitly and actively taking this factor into account.
Motivated by the above observation, we propose a novel Adaptive Language-guided Multimodal Transformer (ALMT) to improve the performance of MSA by addressing the adverse effects of disruptive information in visual and audio modalities. In ALMT, each modality is first transformed into a unified form by using a Transformer with initialized tokens. This operation not only suppresses the redundant information across modalities, but also compresses the length of long sequences to facilitate efficient model computation. Then, we introduce an Adaptive Hyper-modality Learning (AHL) module that uses different scales of language features with dominance to guide the visual and audio modalities to produce the intermediate hyper-modality token, which contains less sentiment-irrelevant information. Finally, we apply a cross-modality fusion Transformer with language features serving as query and hyper-modality features serving as key and value. In this sense, the complementary relations between language and visual and audio modalities are implicitly reasoned, achieving robust and accurate sentiment predictions. In summary, the major contributions of our work can be summarized as:
@@
  • ALMT achieves state-of-the-art performance in several public and widely adopted datasets. We further provide in-depth analysis with rich empirical results to demonstrate the validity and necessity of the proposed approach.
Figure 2: Processing pipeline of the proposed ALMT for multimodal sentiment analysis (MSA). With the multimodal input, we first apply three Transformer layers to embed modality features with low redundancy. Then, we employ a Hyper-modality Learning (AHL) module to learn a hyper-modality representation from visual and audio modalities under the guidance of language features at different scales. Finally, a Cross-modality Fusion Transformer is applied to incorporate hyper-modality features based on their relations to the language features, thus obtaining a complementary and joint representation for MSA.
Figure 3: An example of the Adaptive Hyper-modality Learning (AHL) Layer.
@@ 2.1 Multimodal Sentiment Analysis
As mentioned in the section above, most previous MSA methods are mainly classified into two categories: representation learning-centered methods and multimodal fusion-centered methods.
For representation learning-centered methods, Hazarika et al. (2020) and Yang et al. (2022) argued representation learning of multiple modalities as a domain adaptation task. They respectively used metric learning and adversarial learning to learn the modality-invariant and modality-specific subspaces for multimodal fusion, achieving advanced performance in several popular datasets. Han et al. (2021) proposed a framework named MMIM that improves multimodal fusion with hierarchical mutual information maximization. Rahman et al. (2020) and Guo et al. (2022) devised different architectures to enhance language representation by incorporating multimodal interactions between language and non-verbal behavior information. However, these methods do not pay enough attention to sentiment-irrelevant redundant information that is more likely to be present in visual and audio modalities, which limits the performance of MSA.
For multimodal fusion-centered methods, Zadeh et al. (2017) proposed a fusion method (TFN) using a tensor fusion network to model the relationships between different modalities by computing the cartesian product. Tsai et al. (2019a) and Huang et al. (2020) introduced a multimodal Transformer to align the sequences and model long-range dependencies between elements across modalities. However, these methods directly fuse information from uni-modalities, which is more accessible to the introduction of sentiment-irrelevant information, thus obtaining sub-optimal results.
@@ 3.3 Modality Embedding
$H^{1}_{m}=\operatorname{E^{0}_{m}}(\operatorname{concat}(H^{0}_{m},U_{m}),\theta_{E^{0}_{m}})\in\mathbb{R}^{T\times d}$ (1)
where $H^{1}_{m}$ is the unified feature of each modality $m$ with a size of $T\times d$, $\operatorname{E^{0}_{m}}$ and $\theta_{E^{0}_{m}}$ respectively represent the modality feature extractor and corresponding parameters, $\operatorname{concat}(\cdot)$ represent the concatenation operation.
In practice, $T$ and $d$ are set to 8 and 128, respectively. The structure of the transformer layer is designed as the same as the Vision Transformer (VIT) (Dosovitskiy et al., 2021) with a depth setting of 1. Moreover, it is worth noting that transferring the essential modality information to initialized low-dimensional tokens is beneficial to decrease the redundant information that is irrelevant to human sentiment, thus achieving higher efficiency with lesser parameters.
@@ 3.4.1 Construction of Two-scale Language Features
$H^{i}_{l}=\operatorname{E}^{i}_{l}(H^{i-1}_{l},\theta_{E_{l}^{i}})\in\mathbb{R}^{T\times d}$ (2)
@@ 3.4.2 Adaptive Hyper-modality Learning Layer
where $\operatorname{softmax}$ represents weight normalization operation, $W_{Q_{l}}\in\mathbb{R}^{d\times d_{k}}$ and $W_{K_{a}}\in\mathbb{R}^{d\times d_{k}}$ are learnable parameters, $d_{k}$ is the dimension of each attention head. In practice, we used 8-head attention and set $d_{k}$ to 16.
Similar to $\alpha$, $\beta$ represents the similarity matrix between language modality and visual modality:
@@
where $W_{K_{v}}\in\mathbb{R}^{d\times d_{k}}$ is learnable.
Then the hyper-modality features $H^{j}_{hyper}$ can be updated by weighted audio features and weighted visual features as:
@@ 3.6 Overall Learning Objectives
where $N_{b}$ is the number of samples in the training set, $y^{n}$ is the sentiment label of the $n$-th sample. $\hat{y}^{n}$ is the prediction of our ALMT.
In addition, thanks to our simple optimization goal, compared with advanced methods Hazarika et al. (2020); Yu et al. (2021) with multiple optimization goals, ALMT is much easier to train without tuning extra hyper-parameters. More details are shown in section 4.5.10.
Table 1: Comparison on MOSI and MOSEI. Note: the best result is highlighted in bold; * represents the result is from Hazarika et al. (2020); $\dagger$ represents the result is from Mao et al. (2022) and its corresponding GitHub page4.1.
@@
Table 2: Comparison results on CH-SIMS. Note: the best result is highlighted in bold; $\dagger$ represents the result is from Mao et al. (2022) and its corresponding GitHub page4.1.
@@ 4.1 Datasets
We conducted extensive experiments on three popular trimodal datasets (i.e., MOSI Zadeh et al. (2016), MOSEI (Zadeh et al., 2018), and CH-SIMS (Yu et al., 2020)).
MOSI. The dataset comprises 2,199 multimodal samples encompassing visual, audio, and language modalities. Specifically, the training set consists of 1,284 samples, the validation set contains 229 samples, and the test set encompasses 686 samples. Each individual sample is assigned a sentiment score ranging from -3 (indicating strongly negative) to 3 (indicating strongly positive).
MOSEI. The dataset comprises 22,856 video clips collected from YouTube with a diverse factors (e.g., spontaneous expressions, head poses, occlusions, illuminations). This dataset has been categorized into 16,326 training instances, 1,871 validation instances, and 4,659 test instances. Each instance is meticulously labeled with a sentiment score ranging from -3 to 3. And the sentiment scores from -3 to 3 indicate most negative to most positive.
CH-SIMS. It is a Chinese multimodal sentiment dataset that comprises 2,281 video clips collected from variuous sources, such as different movies and TV serials with spontaneous expressions, various head poses, etc. It is divided into 1,368 training samples, 456 validation samples, and 457 test samples. Each sample is manually annotated with a sentiment score from -1 (strongly negative) to 1 (strongly positive).
11footnotetext: https://github.com/thuiar/MMSA/blob/master/results/result-stat.md
@@ 4.4 Performance Comparison
Table 1 and Table 2 list the comparison results of our proposed method and state-of-the-art methods on the MOSI, MOSEI, and CH-SIMS, respectively.
As shown in the Table 1, the proposed ALMT obtained state-of-the-art performance in almost all metrics. On the task of more difficult and fine-grained sentiment classification (Acc-7), our model achieves remarkable improvements. For example, on the MOSI dataset, ALMT achieved a relative improvement of 1.69% compared to the second-best result obtained by CHFN. It demonstrates that eliminating the redundancy of auxiliary modalities is essential for effective MSA.
Moreover, it is worth noting that the scenarios in SIMS are more complex than MOSI and MOSEI. Therefore, it is more challenging to model the multimodal data. However, as shown in the Table 2, ALMT achieved state-of-the-art performance in all metrics compared to the sub-optimal approach. For example, compared to Self-MM, it achieved relative improvements with 1.44% on Acc-2 and 1.40% on the corresponding F1, respectively. Achieving such superior performance on SIMS with more complex scenarios demonstrates ALMT’s ability to extract effective sentiment information from various scenarios.
@@ 4.5.1 Effects of Different Modalities
To better understand the influence of each modality in the proposed ALMT, Table 3 reports the ablation results of the subtraction of each modality to the ALMT on the MOSI and CH-SIMS datasets, respectively. It is shown that, if the AHL is removed based on the subtraction of each modality, the performance decreases significantly in all metrics. This phenomenon demonstrates that AHL is beneficial in reducing the sentiment-irrelevant redundancy of visual and audio modalities, thus improving the robustness of MSA.
In addition, we note that after removing the video and audio inputs, the performance of ALMT remains relatively high. Therefore, in the MSA task, we argue that eliminating the sentiment-irrelevant information that appears in auxiliary modalities (i.e., visual and audio modalities) and improving the contribution of auxiliary modalities in performance should be paid more attention to.
Table 3: Effects of different modalities. Note: the best result is highlighted in bold.
@@ 4.5.2 Effects of Different Components
Table 4: Effects of different components. Note: the best result is highlighted in bold.
@@
4.5.3 Effects of Different Query, Key, and Value Settings in Fusion Transformer
Table 5: Effect of different Query, Key, and Value settings in Fusion Transformer.
Q K & V MOSI CH-SIMS
Acc-7 MAE Acc-5 MAE
$H_{hyper}^{3}$ $H_{l}^{3}$ 48.10 0.707 44.64 0.410
$H_{l}^{3}$ $H_{hyper}^{3}$ 49.42 0.683 45.73 0.404
@@ 4.5.4 Effects of the Guidance of Different Language Features in AHL
To discuss the effect of the guidance of different language features in AHL, we show the ablation result of different guidance settings on MOSI and CH-SIMS in Table 6. In practice, we replace the AHL layer that do not require language guidance with MLP layer. Obviously, we can see that the ALMT can obtain the best performance when all scals of language features (i.e., $H_{l}^{1}$, $H_{l}^{2}$, $H_{l}^{3}$) involve the guidance of hyper-modality learning.
In addition, we found that the model is more difficult to converge when AHL is removed. It indicates that sentiment-irrelevant and conflicting information visual and audio modalities may limit the improvement of the model.
Table 6: Effects of different guidance of different language features in AHL. Note: the best result is highlighted in bold.
@@
4.5.5 Effects of Different Fusion Techniques
To analyze the effects of different fusion techniques, we conducted some experiments, whose results are shown in the table 7. Obviously, on the MOSI dataset, the use of our Cross-modality Fusion Transformer to fuse language features and hyper-modality features is the most effective. On the CH-SIMS dataset, although TFN achieves better performance on the MAE metric, its Acc-5 is lower. Overall, using Transformer for feature fusion is an effective way.
Table 7: Effects of different fusion techniques.
Fusion Technique MOSI CH-SIMS